OpenStudy (anonymous):

Need help with figuring factor quadratice for 14x^2 + 23x +8

OpenStudy (raden):

14x^2 + 23x + 8 = (14x^2 + 7x) + (16x + 8) = 7x(2x +1) + 8(2x + 1) = (7x + 8)(2x + 1)

OpenStudy (raden):

hoped that make sense

OpenStudy (anonymous):

thank you very much

OpenStudy (raden):

you're welcome

OpenStudy (anonymous):

This is what it is wanting me to do and I can not get the numbers to match up no matter how I do them.